OpenStudy (anonymous):

Evaluate sin 180. Round your answers to the nearest hundredth if necessary.

OpenStudy (solomonzelman):

you can get an exact value though... why do you want to find an approximation ?

OpenStudy (solomonzelman):

sin(180)=sin(90+90)=sin(90)cos(90)+cos(90)sin(90)=2cos(90)sin(90) sin(90)=1 cos(90)=0 Hence 2cos(90)sin(90)=2(0)(1)=0

OpenStudy (anonymous):

 This may help you as well if your a visual person, long as you know your trig ratios :)